# Compactor: Calibrated Query-Agnostic KV Cache Compression with Approximate Leverage Scores

### v1

@@ Abstract
Modern Large Language Models (LLMs) are increasingly trained to support very large context windows. Unfortunately the ability to use long contexts in generation is complicated by the large memory requirement of the KV cache, which scales linearly with the context length. This memory footprint is often the dominant resource bottleneck in real-world deployments, limiting throughput and increasing serving cost. One way to address this is by compressing the KV cache, which can be done either with knowledge of the question being asked (query-aware) or without knowledge of the query (query-agnostic). We present Compactor , a parameter-free, query-agnostic KV compression strategy that uses approximate leverage scores to determine token importance. We show that Compactor can achieve the same performance as competing methods while retaining $\nicefrac{{1}}{{2}}$ the tokens in both synthetic and real-world context tasks, with minimal computational overhead. We further introduce a procedure for context-calibrated compression, which allows one to infer the maximum compression ratio a given context can support. Using context-calibrated compression, we show that Compactor achieves full KV performance on Longbench while reducing the KV memory burden by 63%, on average. To demonstrate the efficacy and generalizability of our approach, we apply Compactor to 27 synthetic and real-world tasks from RULER and Longbench, with models from both the Qwen 2.5 and Llama 3.1 families.
@@ 1 Introduction
BaseRandomSnapKVPyramidKVCompactor 0×0$\times$0 ×1×1$\times$1 ×2×2$\times$2 ×3×3$\times$3 ×Compression RateLlama 3.1Qwen 2.5
Figure 1: Eviction methods with context-calibrated compression on Longbench. Each maintains similar performance as 100% KV, but at different compression levels.
Much effort has been devoted to increasing the context length supported by modern Large Language Models (LLMs). While many LLMs now support incredibly long context windows, practical use of long documents is complicated by the autoregressive attention mechanism: at each generation step we must retain all key/value pairs for every past token, so the KV cache grows linearly with context length. For example, running Qwen-2.5 32B on a prompt of size 100K requires maintenance of 26 GB of KV cache, per request; such a memory burden limits the scalability of long-context deployment. This memory burden is even more pressing, given the widespread use of prompt caching to accelerate inference [32, 1, 28].
To address this, two orthogonal lines of work have emerged. One has focused on improving LLM generation efficiency by careful management of the KV cache [22, 41], and designing more complex prompt caching strategies [15]. The other has focused on methods that reduce the memory burden of the KV cache by compressing, evicting, or summarizing tokens. Less work has focused on the integration of the aforementioned techniques.
We identify two key barriers to the integration of KV compression techniques into efficient LLM serving systems. Firstly, most compression schemes rely on knowing the query at compression time; in this "query-aware" setting, the query is used to inform which tokens are evicted. Unfortunately, the performance of query-aware compression methods degrades significantly when the query is not known at compression time (the "query-agnostic" regime) [7, 25]. Additionally, query-aware schemes are incompatible with multi-query prefix caching: each new question would evict different prefix tokens, so the compressed prefix cannot be reused across prompts. Second, even when query-agnostic compression is possible, determining how many tokens can be evicted without degrading quality is non-trivial. Such determinations are especially needed in applications that lack a prescribed token or memory budget, but where generation quality is paramount. To the best of our knowledge, there has been little prior work targeting this issue of ensuring context-calibrated compression. Our work is orthogonal, and can be composed with, head-level eviction policies as well [35, 33].
@@
In this work we present Compactor , a lightweight, parameter-free KV-eviction strategy that operates without query knowledge and still preserves answer quality. Our contributions are as follows:
  1. (1)
    We propose Compactor , a two-component eviction mechanism that uses leverage scores and non-causal attention scores to significantly outperform current query-agnostic mechanisms. We further propose a fast randomized algorithm to approximate the aforementioned leverage scores to provide a fast, practical implementation of Compactor .
  2. (2)
@@
  3. (3)
    We evaluate Compactor across multiple LLMs, and both synthetic and real-world long-context datasets. We show that on Longbench Compactor performs comparably to using the full KV cache, while halving the memory burden of the KV cache, with minimal computational overhead.
@@ 2 Background
We focus on a Transformer based language model [34]. Applying the standard self-attention mechanism to a context of length $N$ (“prefilling”) has time complexity $O(N^{2})$, while auto-regressively decoding a single token costs $O(N)$; storing the resulting key–value (KV) cache requires $O(N)$ memory. Prefilling is therefore both compute- and memory-bound, whereas decoding is dominated by memory bandwidth. A large body of work tackles these bottlenecks by manipulating the KV cache at different points in the generation pipeline. To describe this work, we adopt the taxonomy of Li et al. [25], which divides KV cache management into 4 stages: KV generation, KV retrieval, KV loading, and KV compression.
@@
KV Generation   The aim at this stage is to construct the cache more cheaply during prefilling. Recent work pursues attempts to sparsify attention patterns during prefilling[4, 36, 20]; these works largely differ in the particular sparsity mask they employ. Another line of work tries to replace self-attention altogether with recurrent models that have bounded state sizes [16, 26]. Yet another line of work focuses on cache-aware scheduling during the prefilling stage [13, 40, 9]. Finally, prompt-level compression can be applied before the model to drop unimportant tokens, lowering prefill cost proportionally [29]. Such architectural changes are usually query-agnostic.
@@
KV Retrieval & Loading  Once a cache exists, the system must decide which cached segments to reuse across requests. Efficient inference frameworks already exploit exact prefix matching[22, 41]. When the desired KVs reside in slower memory, they must be loaded to GPU efficiently. For example, PagedAttention in vLLM treats GPU RAM as a pool of fixed-size pages, enabling fine-grained sharing of KV segments across requests. Similarly, [31] separate the inference workload into prefilling and a decoding cluster, and use a disaggregated KV store to improve (CPU) RAM and SSD utilization. Beyond exact reuse, semantic retrieval matches new prompts to a vector database of cached prefixes, by either skipping the prefill when the prefix similarity exceeds a learned threshold [37], or anticipates which KV elements will be attended to and pre-fetching them [23]. Such methods are inherently query-aware.
@@
KV Compression   The final line of defense is to reduce the footprint of the cache. Quantization is the most direct route, and reduces the memory footprint of each element in the KV cache [27, 38]. Orthogonal works prune tokens altogether to reduce the sequence dimension. StreamingLLM keeps only a sliding window of recent tokens [36]; $\text{H}_{2}\text{O}$, SnapKV, and PyramidKV use accumulated attention scores to drop the least-attended tokens (see §3.3 for further details) [39, 24, 6]. Variants of the above methods apply different compression ratios or strategies to different heads [33]. Token-dropping techniques work well in the query-aware setting, because the attention that tokens in the question place on tokens in the context are used to select necessary tokens. However, in the query-agnostic regime, their performance suffers greatly [25, 7].
@@
Each of these techniques improves KV-cache efficiency along an orthogonal dimension and can therefore be composed to yield highly efficient inference procedures. In this work we focus on KV-cache compression along the sequence dimension; accordingly, we compare only against methods that explicitly reduce the KV cache size via token-dropping in our experiments.
@@ 3 Compactor
Preliminaries & Overview   Let a sequence consist of $N$ tokens, each mapped to $d$-dimensional key and value vectors collected in matrices $\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times d}$. Standard self-attention caches the entire $(\mathbf{K},\mathbf{V})$ matrices, incurring $O(Nd)$ memory burden. We aim to retain only a fraction $r\in(0,1]$ of the tokens while preserving a model’s predictive quality. While not an asymptotic improvement in the memory usage, in practice one can achieve significant memory savings with minimal performance loss by utilizing token eviction. To determine which tokens are to be evicted/discarded, we compute a vector $\vec{s}\in\mathbb{R}^{N}$ of "importance scores" and select the top-$k$ with $k=\lceil r\cdot N\rceil$ highest scoring tokens for retention. To compute $\vec{s}$, we introduce two complementary score families: geometry-based outlier scores $\vec{o}$ and task-driven attention scores $\vec{a}$; each captures different, essential, aspects of the key distribution. We blend them to generate final importance scores $\vec{s}$.
Notation  We consider a transformer-based language model [34], denoted by LM , over a vocabulary $\mathcal{V}$. Let $c\in\mathcal{V}^{N}$ denote a context, and with slight abuse of notation, let $\tilde{c}_{r}$ denote (the KV cache of) a context that has been compressed with retention rate $r$ (i.e if $r=0.1$, we retain 10% of tokens). Throughout, indices $i,j\in[N]$ refer to tokens, whereas boldface capitals denote matrices. For a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, $A_{i}$ is its $i$-th row, $\operatorname{rank}(\mathbf{A})$ its rank, and $\mathbf{A}^{\top}$ its transpose. We denote the total, outlier, and attention "importance score" vectors as $\vec{s},\vec{o},\vec{a}$ respectively. Let $\bar{{v}}$ denote the average value of a vector $\vec{{v}}$, and $\operatorname{std}(\vec{{v}})$ denote the standard deviation of the values of the vector. For clarity, we do not notate the head dimension, as the same procedure is applied to heads independently.
@@ 3.1 Outlier Scores via Statistical Leverage
We first turn our attention to generating “outlier" scores $\vec{o}\in\mathbb{R}^{N}$. Retaining merely the most-attended keys fails when a key is rarely queried yet encodes unique information, i.e., it is an outlier in the key space. Eviction of these outliers leads to irreversible information loss. We formalize the notion of "outlierness" via the classical concept of statistical leverage:
@@
$\ell_{i}=||U_{i}||_{2}^{2}\qquad\qquad\sum_{i=1}^{N}\ell_{i}=\operatorname{rank}(k)$
Intuitively, rows whose associated left-singular vectors have large Euclidean norm align closely with the directions in which $\mathbf{K}$ has the greatest variance. Omitting these rows would discard significant information. Theorem 1 formalizes the utility of leverage scores in row selection.
@@
Let $\epsilon,\delta\in(0,1)$ and data matrix $\mathbf{K}\in\mathbb{R}^{N\times d}$ be given and take $k\geq Cd\log(\frac{d}{\delta})\epsilon^{-2}$ for some universal constant $C$. Construct $\hat{\mathbf{K}}_{k}\in\mathbb{R}^{k\times d}$ by sampling $k$ times (with replacement) from the distribution that places probability mass $\ell_{i}/d$ on the $i$-th row. Scale each of the $k$ chosen rows by $\sqrt{d/(k\ell_{i})}$. Then, with probability $1-\delta$, we have
@@
where $A\preccurlyeq B$ means that $B-A$ is PSD. In practice [5, 30, 21], even deterministically choosing rows corresponding to the top-$k$ leverage scores provides a good spectral sparsification of $K$.
@@
where $\mathbf{K}=\mathbf{U\Sigma V}^{\top}$ denotes the SVD of pre-rope keys. Naively computing leverage scores by computing the SVD of an $n\times d$ matrix $\mathbf{K}=\mathbf{U\Sigma V}^{\top}$ is much too slow in practice. Instead we utilize the following simple identity:
@@ 3.2 Approximate Leverage Score Computation
Given the computational intensity of computing leverage scores, there has been substantial work in developing methods for approximating them [11, 8, 12, 42]. A key technique in most leverage score approximation algorithms is “sketching” the original matrix to obtain a smaller matrix that roughly preserves the properties of the original, in the following sense:
@@
3:$\mathbf{\Phi}\in\mathbb{R}^{d\times k}$ with $\Phi_{ij}\sim\mathcal{N}(0,\frac{1}{k})$
4:$\hat{\mathbf{K}}\leftarrow\mathbf{K\Phi}$
5:$\mathbf{G}\leftarrow\hat{\mathbf{K}}^{\top}\hat{\mathbf{K}}$
6:$\textsc{svd}(\mathbf{G})=\tilde{\mathbf{V}}{\mathbf{\Sigma}}^{2}\tilde{\mathbf{V}}^{\top}$triangleright SVD of a $k\times k$ matrix
7:$\tilde{\mathbf{U}}\leftarrow\hat{\mathbf{K}}\tilde{\mathbf{V}}\tilde{\mathbf{\Sigma}}^{-1}$
8:return $\vec{o}\leftarrow\operatorname{diag}(\tilde{\mathbf{U}}\tilde{\mathbf{U}}^{\top})$
We refer to $\mathbf{\Phi}$ as a “sketching" matrix. In contrast to prior work in leverage-score approximation, which applies left sketches ($\mathbf{\Phi}\mathbf{K}$), we apply a right sketch ($\mathbf{K}\mathbf{\Phi}$) to reduce the dimensionality of the column-space of the data, at the cost of weaker theoretical bounds on the error in the approximate leverage scores. The reason for using right sketching is three-fold: (1) speeding up computation of Eq 2 requires a right sketch; (2) instantiating a left sketching matrix and subsequently computing $\mathbf{\Phi}\mathbf{K}$ requires substantial memory when $N$ is large; and (3) in practice, we observe no degradation in the utility of leverage scores for KV cache approximation when right sketching (see § 5.5). The algorithm to compute approximate leverage scores is presented in Algorithm 1. The quality of the (Gaussian) right-sketching leverage score approximation is formalized in the following theorem:
@@
Let data matrix $\mathbf{K}\in\mathbb{R}^{N\times d}$ and target dimension $k=\lceil r\cdot N\rceil$ be given. Let distortion factor $\epsilon\in(0,1)$ and failure probability $\delta\in(0,1)$ be given. Define $\mathbf{\Phi}\in\mathbb{R}^{d\times k}$ be a matrix whose entries are drawn i.i.d from $\mathcal{N}(0,\frac{1}{k})$. Compute the SVD of the sketched matrix $\mathbf{K\Phi}$ and approximate leverage scores ~$\tilde{\ell}_{i}$ as:
$\textsc{svd}\left(\mathbf{K}\mathbf{\Phi}\right)=\mathbf{\tilde{U}}\mathbf{\tilde{\Sigma}}\mathbf{\tilde{V}}^{\top}\qquad\qquad\tilde{\ell}_{i}=||\tilde{U}_{i}||_{2}^{2}$
@@
The proof of the theorem is deferred to Appendix A. Other choices of sketching matrix are also admissible. In Appendix D, we show that choosing $\Phi$ to be a Subsampled Randomized Hadamard Transform doesn’t impact downstream performance, further demonstrating the robustness of the technique. Furthermore, note that the particular construction of ~$\tilde{\ell}_{i}$ shown in Algorithm 1 is not the only one possible; one can also use an economy QR decomposition, or eigendecomposition. We find these also work well in practice. Appendix C discusses other methods for computing approximate leverage scores, which may be more suitable for other hardware. In summary, we apply Algorithm 1 to the (pre-position embedded) key-embeddings to generate outlier scores $\vec{o}$. With (approximate) outlier scores in hand, we next derive complementary attention-based scores.

### v2

# Compactor: Calibrated KV Cache Compression with Approximate Leverage Scores

@@ Abstract
Modern Large Language Models (LLMs) are increasingly trained to support very large context windows. Unfortunately the ability to use long contexts in generation is complicated by the large memory requirement of the KV cache, which scales linearly with the context length. This memory footprint is often the dominant resource bottleneck in real-world deployments, limiting throughput and increasing serving costs. One way to address this is by compressing the KV cache, which can be done either with knowledge of the question being asked (query-aware) or without knowledge of the query (query-agnostic). We present Compactor , a training-free, query-agnostic KV compression strategy that uses approximate leverage scores to determine token importance. We show that Compactor can achieve the same performance as competing methods while retaining 20% fewer tokens in both synthetic and real-world context tasks, while being far more task-robust. We further introduce a procedure for context-calibrated compression: inferring the maximum compression a given context supports before significant performance loss. Using context-calibrated compression, we show that Compactor achieves full KV performance on Longbench while reducing the KV memory burden by 68%, on average. To demonstrate the efficacy and generalizability of our approach, we apply Compactor to 27 synthetic and real-world tasks from RULER and Longbench, with models from both the Qwen 2.5 and Llama 3.1 families. Finally, we release compactor-vllm [^1] Our code can be found at https://github.com/vnchari/compactor-vllm., an inference engine and suite of optimized Triton kernels designed to efficiently support the sparse, non-contiguous memory access patterns inherent to compressed KV caches. This work demonstrates that Compactor offers a practical, high-performance solution for alleviating the memory bottleneck in modern LLM deployment.
@@ 1 Introduction
Much effort has been devoted to increasing the context length supported by modern Large Language Models (LLMs). While many LLMs now support incredibly long context windows, practical use of long documents is complicated by the autoregressive attention mechanism: at each generation step we must retain all key/value pairs for every past token, so the KV cache grows linearly with context length. For example, running Qwen-2.5 32B on a prompt of size 100K requires maintenance of 26 GB of KV cache, per request; such a memory burden limits the scalability of long-context deployment. This memory burden is especially problematatic for inference servers, given the widespread use of prompt caching.
To address this, two orthogonal lines of work have emerged. One has focused on improving LLM generation efficiency by careful management of the KV cache (kwon_efficient_2023; zheng_sglang_2024), and designing more complex prompt caching strategies (gim_prompt_2024). The other has focused on methods that reduce the memory burden of the KV cache by compressing, evicting, or summarizing tokens. Less work has focused on the integration of the aforementioned techniques.
We identify two key barriers to the integration of KV compression techniques into efficient LLM serving systems. Firstly, most compression schemes rely on knowing the query at compression time; in this query-aware setting, the query is used to inform which tokens are evicted. Unfortunately, the performance of query-aware compression methods degrades significantly when the query is not known at compression time (the query-agnostic regime) (chari_kv-distill_2025; li_scbench_2025). Additionally, query‐aware schemes are incompatible with multi‐query prefix caching: each new question would evict different prefix tokens, so the compressed prefix cannot be reused across prompts. Second, even when query-agnostic compression is possible, determining how many tokens can be evicted without degrading quality is non-trivial. Such determinations are especially needed in applications that lack a prescribed token or memory budget, but where generation quality is paramount. To the best of our knowledge, there has been little prior work targeting this issue of context-calibrated compression.
@@
In this work we present Compactor , a lightweight, (trainable) training-free KV-eviction strategy that operates without query knowledge and still preserves answer quality. Our contributions are as follows:
  1. 1.
    We propose Compactor , a two-component eviction mechanism that uses leverage scores and non-causal attention scores to significantly outperform current query-agnostic mechanisms. We further propose a fast randomized algorithm to approximate the aforementioned leverage scores to provide a fast, practical implementation of Compactor , and prove guarantees regarding the quality of the output of the randomized algorithm.
  2. 2.
@@
  3. 3.
    We evaluate Compactor across multiple LLMs, and both synthetic and real-world long-context datasets. We show that on Longbench, Compactor performs comparably to using the full KV cache, while retaining a third of the KV cache. On synthetic UUID needle retrieval tasks, we show that we can achieve full performance while retaining only 50% of the cache.
  4. 4.
    We release compactor-vllm, an inference engine with a custom paged KV cache implementation and a suite of custom Triton kernels to make KV cache compression practically feasible.
@@ 2 Background
We focus on a Transformer based language model (vaswani_attention_2023). Applying the standard self‑attention mechanism to a context of length $N$ (“prefilling”) has time complexity $O(N^{2})$, while auto‑regressively decoding a single token costs $O(N)$; storing the resulting key–value (KV) cache requires $O(N)$ memory. Prefilling is therefore usually compute‑bound, whereas decoding is dominated by memory bandwidth. A large body of work tackles these bottlenecks by manipulating the KV cache at different points in the generation pipeline. To describe this work, we adopt the taxonomy of li_scbench_2025, which divides KV cache management into 4 stages: KV generation, KV retrieval, KV loading, and KV compression.
@@
KV Generation   The aim at this stage is to construct the cache more cheaply during prefilling. Recent work pursues attempts to sparsify attention patterns during prefilling(beltagy_longformer_2020; xiao_efficient_2024; jiang_minference_2024); these works largely differ in the particular sparsity mask they employ. Another line of work tries to replace self‑attention altogether with recurrent models that have bounded state sizes (gu_mamba_2024; lieber_jamba_2024). Yet another line of work focuses on cache-aware scheduling during the prefilling stage (fu_lazyllm_2024; zhao_alisa_2024; cho_kv-runahead_2024). Finally, prompt‑level compression can be applied before the model to drop unimportant tokens, lowering prefill cost proportionally (pan_llmlingua-2_2024). Such architectural changes are usually query-agnostic.
@@
KV Retrieval & Loading  Once a cache exists, the system must decide which cached segments to reuse across requests. Efficient inference frameworks already exploit exact prefix matching(kwon_efficient_2023; zheng_sglang_2024). When the desired KVs reside in slower memory, they must be loaded to GPU efficiently. For example, PagedAttention in vLLM treats GPU RAM as a pool of fixed‑size pages, enabling fine‑grained sharing of KV segments across requests. Similarly, (qin_mooncake_2024) separate the inference workload into prefilling and a decoding cluster, and use a disaggregated KV store to improve (CPU) RAM and SSD utilization. Beyond exact reuse, semantic retrieval matches new prompts to a vector database of cached prefixes, by either skipping the prefill when the prefix similarity exceeds a learned threshold (yao_cacheblend_2024), or anticipates which KV elements will be attended to and pre-fetching them (lee_infinigen_2024). Such methods are inherently query-aware.
@@
KV Compression   The final line of defense is to reduce the footprint of the cache. Quantization is the most direct route, and reduces the memory footprint of each element in the KV cache (liu_kivi_2023; zhang_kv_2024). Orthogonal works prune tokens altogether to reduce the sequence dimension. StreamingLLM keeps only a sliding window of recent tokens (xiao_efficient_2024); $\text{H}_{2}\text{O}$, SnapKV, and PyramidKV use accumulated attention scores to drop the least‑attended tokens (see §3.3 for further details) (zhang_h_2o_2023; li_snapkv_2024; cai_pyramidkv_2024). Variants of the above methods apply different compression ratios or strategies to different heads (tang_razorattention_2024; feng_ada-kv_2025). Token-dropping techniques work well in the query-aware setting, because the attention that tokens in the question place on tokens in the context are used to select necessary tokens. However, in the query-agnostic regime, their performance suffers greatly (li_scbench_2025; chari_kv-distill_2025).
@@
Each of these techniques improves KV‑cache efficiency along an orthogonal dimension and can therefore be composed to yield highly efficient inference procedures. In this work we focus on KV‑cache compression along the sequence dimension; accordingly, we compare only against methods that explicitly reduce the KV cache size via token-dropping in our experiments.
@@ 3 Compactor
Preliminaries & Overview   Let a sequence consist of $N$ tokens, each mapped to dd‑dimensional key and value vectors collected in matrices $\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times d}$. Standard self‑attention caches the entire $(\mathbf{K},\mathbf{V})$ matrices, incurring $O(Nd)$ memory burden. We aim to retain only a fraction $r\in(0,1]$ of the tokens while preserving a model’s predictive quality. While not an asymptotic improvement in the memory usage, in practice one can achieve significant memory savings with minimal performance loss by utilizing token eviction. To determine which tokens are to be evicted/discarded, we compute a vector $\vec{s}\in\mathbb{R}^{N}$ of “importance scores” and select the top-$k$ with $k=\lceil r\cdot N\rceil$ highest scoring tokens for retention. To compute $\vec{s}$, we introduce two complementary score families: geometry-based outlier scores $\vec{o}$ and task-driven attention scores $\vec{a}$; each captures different, essential, aspects of the key distribution. We blend them to generate final importance scores $\vec{s}$.
Notation  We consider a transformer-based language model (vaswani_attention_2023), denoted by LM , over a vocabulary $\mathcal{V}$. Let $c\in\mathcal{V}^{N}$ denote a context, and with slight abuse of notation, let $\tilde{c}_{r}$ denote (the KV cache of) a context that has been compressed with retention rate $r$ (i.e if $r=0.1$, we retain 10% of tokens). Throughout, indices $i,j\in[N]$ refer to tokens, whereas boldface capitals denote matrices. For a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, $A_{i}$ is its ii‑th row, $\operatorname{rank}(\mathbf{A})$ its rank, and $\mathbf{A}^{\top}$ its transpose. We denote the total, outlier, and attention “importance score” vectors as $\vec{s},\vec{o},\vec{a}$ respectively. Let $\bar{{v}}$ denote the average value of a vector $\vec{{v}}$, and $\operatorname{std}(\vec{{v}})$ denote the standard deviation of the values of the vector. For clarity, we do not notate the head dimension, as the same procedure is applied to heads independently.
@@ 3.1 Outlier Scores via Statistical Leverage
We first turn our attention to generating “outlier” scores $\vec{o}\in\mathbb{R}^{N}$. Retaining merely the most‐attended keys fails when a key is rarely queried yet encodes unique information, i.e., it is an outlier in the key space. Eviction of these outliers leads to irreversible information loss. We formalize the notion of “outlierness” via the classical concept of statistical leverage:
@@
[…]
Intuitively, rows whose associated left‐singular vectors have large Euclidean norm align closely with the directions in which $\mathbf{K}$ has the greatest variance. Omitting these rows would discard significant information. Theorem 1 formalizes the utility of leverage scores in row selection.
@@
Let $\epsilon,\delta\in(0,1)$ and data matrix $\mathbf{K}\in\mathbb{R}^{N\times d}$ be given and take $k\geq Cd\log(\frac{d}{\delta})\epsilon^{-2}$ for some universal constant $C$. Construct $\hat{\mathbf{K}}_{k}\in\mathbb{R}^{k\times d}$ by sampling $k$ times (with replacement) from the distribution that places probability mass $\tfrac{\ell_{i}}{\operatorname{rank}(\mathbf{K})}$ on the $i$-th row. Scale each of the $k$ chosen rows by $\sqrt{d/(k\ell_{i})}$. Then, with probability $1-\delta$, we have
@@
where $A\preccurlyeq B$ means that BAB-A is PSD. In practice (broadbent_subset_2010; paschou_pca-correlated_2007; jolliffe_discarding_1972), even deterministically choosing rows corresponding to the top-$k$ leverage scores provides a good spectral sparsification of $K$.
@@
where $\mathbf{K}=\mathbf{U\Sigma V}^{\top}$ denotes the SVD of pre-positionally encoded keys. Naively computing leverage scores by computing the SVD of an $n\times d$ matrix $\mathbf{K}=\mathbf{U\Sigma V}^{\top}$ is much too slow in practice. Instead we utilize the following simple identity:
@@ 3.2 Approximate Leverage Score Computation
Given the computational intensity of computing leverage scores, there has been substantial work in developing methods for approximating them (drineas_fast_2012; chen_fast_2021; eshragh_salsa_2023; zuo_quantum-inspired_2021). A key technique in most leverage score approximation algorithms is “sketching” the original matrix to obtain a smaller matrix that roughly preserves the properties of the original, in the following sense:
@@
We refer to $\mathbf{\Phi}$ as a “sketching" matrix. In contrast to prior work in leverage-score approximation, which applies left sketches ($\mathbf{\Phi}\mathbf{K}$), we apply a right sketch ($\mathbf{K}\mathbf{\Phi}$) to reduce the dimensionality of the column-space of the data, at the cost of weaker theoretical bounds on the error in the approximate leverage scores. The reason for using right sketching is three-fold: (1) speeding up computation of Eq 2 requires a right sketch; (2) instantiating a left sketching matrix and subsequently computing $\mathbf{\Phi}\mathbf{K}$ requires substantial memory when $N$ is large; and (3) in practice, we observe no degradation in the utility of leverage scores for KV cache approximation when right sketching (see § 5.5). The algorithm to compute approximate leverage scores is presented in Algorithm 1.
@@
[…]
The quality of the (Gaussian) right-sketching leverage score approximation is formalized in the following theorem:
@@
Let data matrix $\mathbf{K}\in\mathbb{R}^{N\times d}$ and target dimension $k$ be given. Let distortion factor $\epsilon\in(0,1)$ and failure probability $\delta\in(0,1)$ be given. Define $\mathbf{\Phi}\in\mathbb{R}^{d\times k}$ be a matrix whose entries are drawn i.i.d from $\mathcal{N}(0,\frac{1}{k})$. Compute the SVD of the sketched matrix $\mathbf{K\Phi}$ and approximate leverage scores $\tilde{\ell}_{i}$ as:
[…]
@@
The proof of the theorem is deferred to Appendix A. Other choices of sketching matrix are also admissible. In Appendix D, we show that choosing $\Phi$ to be a Subsampled Randomized Hadamard Transform doesn’t impact downstream performance, further demonstrating the robustness of the technique. Furthermore, note that the particular construction of $\tilde{\ell}_{i}$ shown in Algorithm 1 is not the only one possible; one can also use an economy QR decomposition, or eigendecomposition. We find these also work well in practice. Appendix C discusses other methods for computing approximate leverage scores, which may be more suitable for other hardware or for highly ill-conditioned key matrices. In summary, we apply Algorithm 1 to the (pre-position embedded) key-embeddings to generate outlier scores $\vec{o}$. With (approximate) outlier scores in hand, we next derive complementary attention-based scores.